OpenStudy (anonymous):

Pies are ordered from the Pizza Palace according to their diameter. Large:sixteen inches Medium:fourteen inches Small:twelve inches How many more square inches are in the large pizza than in the small pizza? twelve point five six square inches forty-seven point one zero square inches eighty-seven point nine two square inches three hundred fifty-one point six eight square inches

OpenStudy (bradely):

Large-small =pi/4(16^2-12^2) =pi/4(256-144) =112pi =351.68 square inches So option 4 is correct Source: http://www.mathskey.com/question2answer/